# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v2

@@ 3 Problem Formulation
8:  Average neighbors: $\mathbf{X}_{\text{matched}}\leftarrow\text{mean of neighbors for each sample}$
9:  Construct pseudo-pairs: $\mathcal{C}^{\prime}\leftarrow\{(\mathbf{X}_{\text{sample}},\mathbf{X}_{\text{matched}})\}$
10:  Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
11:  Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
12:end for
Algorithm 4 Refine-2: Clustering-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Smoothing parameter $\alpha$, number of clusters $c^{\prime}$
3:Refined transformation matrix $\mathbf{W}$ after clustering-based refinement
4:Cluster space A: $\{\mathbf{c}^{A}_{j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{A},c^{\prime})$
5:Transform centroids: $\{\mathbf{c}^{A\to B}_{j}\}\leftarrow\{\mathbf{c}^{A}_{j}\mathbf{W}\}$
6:Cluster space B with seeds: $\{\mathbf{c}^{B}_{j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{B},c^{\prime},\text{init}=\{\mathbf{c}^{A\to B}_{j}\})$
7:Construct pairs $\mathcal{C}^{\prime}$ by matching cluster centroids
8:Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
9:Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
10:return $\mathbf{W}$
@@
For the gtr-gte pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption. Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
Our linear approach addresses key limitations of existing methods: GANs are notoriously unstable and hard to train with many hyperparameters, require extensive computational resources and training time, and often need multiple seeds to achieve good solutions. In contrast, we know that in supervised cases (model stitching), there usually exists an affine mapping between latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. Our method requires significantly less data (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles Hoshen and Wolf (2018)’s approach for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This scheme is very powerful, and we apply a similar method. We find that the approximate matching stage especially requires more delicate treatment in our case, so their approach is not applicable as-is.
The stages incrementally improve alignment quality – each stage builds upon structural correspondences to establish increasingly accurate alignments:
  • Approximate Matching: Find approximate pairs of embeddings from both spaces using landmark points as anchors for relative representations
  • Mapping Estimation: Learn a linear orthogonal mapping from approximate matchings, where linearity and orthogonality provide an inductive bias to smooth out incorrect pairings
  • Iterative Refinement: Use the initial solution as a stepping stone for better matching/mapping, iterating until convergence or budget exhaustion
The overall process is described in Algorithm 1.
4.2 Approximate Matching
Preprocessing.
We first address arbitrary translations and scalings across embedding spaces. Both spaces are centered around their respective means and normalized to the unit hypersphere:
Centroid Matching.
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. Each run consists of k-means clustering followed by solving a quadratic assignment problem (QAP) for cluster correspondence.
In each run, we first perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute structural similarity matrices between cluster centroids:
We find the optimal matching between cluster centroids by solving the quadratic assignment problem:
$\mathbf{P}^{*}=\operatorname{argmax}_{\mathbf{P}\in\Pi_{k}}\text{Tr}\left(\mathbf{S}_{A}^{T}\mathbf{P}\mathbf{S}_{B}\mathbf{P}^{T}\right)$
where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices. For solving QAP, we use $2$-OPT (Croes, 1958). We run it 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment over a single run of 2-OPT. Due to the small number of cluster centroids, this takes only a few seconds.
Relative Representations.
Once we have identified an alignment between centroids, we use the aligned centroids as anchors. For each run within the anchor discovery phase, we represent each embedding through its relationships to anchor points. For embedding $\mathbf{x}^{A}_{i}$ in a given run, we construct:
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]^{T}$
where $\mathbf{r}^{A}_{i}$ is the relative representation of the $i$-th embedding in space $A$ with respect to the $k$ cluster centroids from that run.
Robustification through Ensembling.
We improve robustness by concatenating relative representations from multiple runs:
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk}$
This concatenation creates richer structural signatures that are more robust to clustering initialization randomness.
Anchor alignment algorithm is provided in Algorithm 2.
4.3 Mapping Estimation
We construct pseudo-parallel pairs by matching embeddings by similarity in relative space, averaging the $k$ nearest neighbors to reduce noise:
$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\mathbf{x}^{B}_{j}\right):i=1,\ldots,n_{A}\right\}$
where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space. Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$. The orthogonal transformation is obtained via Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
4.4 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima. Pseudo-code is provided in Algorithms 3 & 4.
Refine-1: Matching-based Refinement.
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, and averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.3), with the crucial difference that now the similarity is performed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100) using subsampling to boost computational efficiency.
We use exponential smoothing for transformation updates:
In the experiments below, we use $\alpha=0.5$.
Refine-2: Clustering-based Refinement.
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing. [^2] Note that the above justification for Refine-2 is our a priori rationale for using this method, and we do not make any claims as to whether this is the true reason for the ability of this method to improve the alignment. Our only claim is that we observe this step improves the alignment consistently, and despite the fact that the Refine-1 has already converged and no longer shows improvement.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Curiously, two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
5 Experiments
We replicate the in-distribution experiment from Jha et al. (2025). We run the algorithm for all pairs among the 5 sets of text encoders used in Jha et al. (2025). The encoders are summarized in Table 2. Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT.
We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime, and repeated three times for each pair to demonstrate consistency. We compare the results to the numbers reported in Jha et al. (2025).
Metrics.
We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm. Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry. Our setup does not change the geometry. [^3] Apart from translation and scaling, which, formally speaking, are a form of transition to a third space, but in a very restricted sense.
Hyperparameters.
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. We use $k=50$ neighbors in Mapping Estimation. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$.
Note this is a rather economical choice of hyperparameters, which we would not suggest using a priori. It turned out to work well, regardless. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the gtr-gte pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
Results.
@@
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is more noticeable in the rank metric. Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $e5$-gtrgtr pair or the gtrgtr-gtegte pair.
We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise. The superior results are an additional benefit, but even where the results are not superior, but rather matching performance, it is achieved for a fraction of the cost. Moreover, even where vec2vec collapses, we observe mini-vec2vec still reaches good performance. Computationally, one run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
Analysis.
@@
We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
@@
gtr 110M T5 Ni et al. 2021
e5 109M BERT Wang et al. 2022
gte 109M BERT Li et al. 2023
stella 109M BERT Zhang et al. 2023
granite 278M RoBERTa Awasthy et al. 2024
@@ 6 Discussion
The mini-vec2vec algorithm is reasonably stable under different choices of parameters, often converging to similar stationary points. Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and frequently have a monotone behavior. For example, generally speaking, increasing the following hyperparameters should improve the outcome (or at least not harm). Importantly, while some of these trends have been observed empirically in some setups, a full experimental analysis has not been made, and so the explanations below on how to choose hyperparameters are suggestions at best, and should not be trusted blindly.
@@
  • Number of concatenated representations: The more concatenated alignments we use, the more likely it is for noisy assignments to be diluted, emphasizing the signal from true assignments. Note that this helps stabilize the algorithm, meaning it is less likely to meet failure modes in the average and worst cases. However, even with fewer alignments, good alignments can be reached, but the probability of failure modes might increase.
@@
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows. This makes it a safe bet to increase the number of iterations if one is unsure whether convergence has been reached. Across all cases examined, we found convergence to be quite fast. Almost without exception, 100 iterations were more than enough, by a large margin.
@@
  • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise. The larger the sample size is, the closer this approximation is to the data. We generally consider higher sampling rates to be better.
@@
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While it gives us more landmark points for alignment, it can theoretically lead to instabilities, because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, to a point. It helps reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is in contrast to multilingual word embeddings, where essentially every point has a ground-truth translation, and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance eventually. Finally, increasing the exponential smoothing parameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. This is likely due to overfitting. The smoothing parameter is implicitly responsible for the maximum distance of the matrix from being orthogonal. While this (alleged) overfitting phenomenon has been observed by us empirically, we did not explore it rigorously, and it might not be a common case. Additionally, increasing $\alpha$ makes convergence slightly slower.
@@
Unlike many deep learning setups – GANs in particular – the range of hyperparameters that lead to strong performance is generously wide. Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter in order to improve effectiveness, if needed.
Failure Modes

### v3

@@ 4.3 Approximate Matching
Robustification via Ensembling
We improve robustness by repeating the process described above and concatenating relative representations from multiple runs:
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk},\qquad\text{similarly for space }B$
This concatenation creates richer structural signatures that are more robust to noise that comes from bad alignments.
4.4 Mapping Estimation
We construct pseudo-parallel pairs by sending each element in space $A$ to the average of its $k$ nearest neighbors from space $B$ (based on similarity in relative space):
$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\hskip-3.00003pt\sum_{\tiny\hskip 2.94443pt{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\hskip-13.25018pt}\hskip-3.00003pt\mathbf{x}^{B}_{j}\,\,\,\,\right):\,i=1,\ldots,n_{A}\right\}$
where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space. Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$. Optimal orthogonal transformation is obtained by Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
4.5 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. Pseudo-code is provided in Listing LABEL:lst:refine.
Refine-1: Matching-based Refinement
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.4), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100), using subsampling to boost computational efficiency further.
We use exponential smoothing for transformation updates:
Refine-2: Clustering-based Refinement
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
def refine_1(X_A, X_B, W):
for t in range(num_iters):
X_sample = sample(X_A, subsampling_rate)
X_transformed = X_sample @ W
neighbors = knn(X_transformed, X_B, num_neighbors=k2)
X_sample_matched = average_nearest_neighbors(X_B, neighbors)
W_new = procrustes(X_sample, X_sample_matched)
W = (1 - alpha) * W + alpha * W_new
return W
def refine_2(X_A, X_B, W):
C_A = k_means(X_A, num_clusters2)
C_B = k_means(X_B, num_clusters2, init=C_A @ W)
W_new = procrustes(C_A, C_B)
return (1 - alpha) * W + alpha * W_new
Listing 2: Pseudo-code for the iterative refinement step
5 Experiments
We replicate the main experiment from jha2025vec2vec. We run our algorithm for all pairs of encoders used in jha2025vec2vec. Details about the encoders are provided in Table 2. Experiments use k-means clustering from scikit-learn (scikit-learn) and QAP is solved with the scipy (2020SciPy-NMeth) implementation of 2-OPT. We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
We take a random subset of the Natural Questions dataset (natural_questions) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate consistency. We compare our results to the numbers reported in jha2025vec2vec.
Metrics
We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by:
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm. Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry. Our setup does not change the geometry (apart from aparametric translation and scaling in the preprocessing step).
Hyperparameters
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We make an economical choice of hyperparameter values due to the large number of runs, but we find the choice to be sufficiently robust still. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. In Mapping Estimation, we use $k=50$ neighbors. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the $\textit{gtr}\leftrightarrow\textit{gte}$ pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
@@ 5 Experiments
For the $gtr\leftrightarrow gte$ pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
Results
@@
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is most noticeable in the rank metric. Importantly, the method is extremely robust as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $\textit{e5}\leftrightarrow\textit{gtr}$ pair or the $\textit{gtr}\leftrightarrow\textit{gte}$ pair. We underline scores that are within a $0.01$ margin, as they might result from rounding noise, stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is good enough.
Analysis
The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that it is achieved for a fraction of the cost. One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware. Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance.
@@
Analyzing the convergence of our algorithm, we observe that despite variation in the approximate matching step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two methods appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
@@
gtr 110M T5 gtr 2021
e5 109M BERT e5 2022
gte 109M BERT gte 2023
stella 109M BERT stella 2023
granite 278M RoBERTa granite 2024
@@ 6.1 Choice of Hyperparameters
mini-vec2vec is stable under different choices of parameters, often converging to similar stationary points. Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and often have a monotone behavior. Here we share trends that we have observed empirically in some setups, but a rigorous experimental analysis has not been made.
@@
  • Number of concatenated representations: Increasing the number of runs in the anchor discovery step reduces the number of the failure modes. This is because noisy assignments are diluted, emphasizing the signal from true assignments. Note that even with very few runs, good alignments can be reached at random but with smaller probability.
@@
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows, unlike some algorithms where increasing information might cause degradation. This makes it a safe bet to increase the number of iterations. Across all cases examined, though, we found convergence to be quite fast. Almost without exception, 100 iterations were enough, by a large margin.
@@
  • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise. The larger the sample size is, the closer this approximation is to the data. As a rule of thumb, we generally considered higher sampling rates to be better (at the cost of slower performance), but we haven’t tested this rigorously.
@@
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While increasing it will give us more landmark points for alignment, it can theoretically lead to instabilities because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, but it eventually leads to diminishing returns, and might even deteriorate performance when neighborhoods become very large. At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation, and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance. Finally, increasing $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. $\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal. Additionally, increasing $\alpha$ makes convergence slightly slower.
@@
In summary, unlike many deep learning setups, GANs in particular, the range of hyperparameters that lead to strong performance is generously wide. Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter to improve effectiveness, if needed.
6.2 Failure Modes